I have to choose if it's two rational, one rational, 2irrational or imaginary solutions.
1. Sē-5s-6=0
2. Tē+12t+36=0
3. Vē-8v-3=0

Answer by stanbon(75887)

I have to choose if it's two Real, one Real, 2irrational or imaginary solutions.
Use the discriminant: b^2 -4ac
----------------------------------
1. Sē-5s-6=0
5^2 - 4*-6 > 0, so 2 Real Number solutions
------------------------------
2. Tē+12t+36=0
12^2-4*36 = 0, so one Real solution with multiplicity 2
------------------------------
3. Vē-8v-3=0
8^2-4*-3 > 0, so 2 Real Number solutions
======================================
Cheers,
Stan H.